Percentage of normal probability distribution

9. The average price of cell phones manufactured by Ahmadi, Inc. is $98 with a standard deviation of $12. Furthermore, it is known that the prices of the cell phonesmanufactured by Ahmadi are normally distributed.


d. If 7,218 of the Ahmadi cell phones were priced at least $119.00, how many cell phones were produced by Ahmadi, Inc.?
Answer #1
known
normal distribution
µ = 98
s = 12
a. % with price of at least 120
z = (x - µ)/s
z = (120 - 98)/12
z = 1.833
want P(x <= 120) = P(z <= 1.8333) = 0.5+0.4664 = 0.9664
so 96.6 %

b. % with a price of at least 81.8
z = (81.8 - 98)/12 = -1.35
want P(x <= 81.80) = P(z<=-1.35) = 0.0885
so 8.85%

c. min and max of middle 95%
area is 0.9500
.9500/2 = 0.475
so
z = -1.96
and
z = +1.96
x = sz + µ
x = (12)(-1.96) + 98
x = $74.48
and
z=1.96
x = (12)(1.96) + 98
x = $121.52
so the min and max prices of the middle 95% are
$74.48 and $121.52
answered by: Danny!!! =]
Answer #2

jhbxfgchgvjbknlm;,';.'?
Answer #3

suck it .. my a$$